Sol Summary – February 15th

Crew 292 Sol Summary Report 15-02-2024

Sol: Sol 11

Summary Title: Our rock samples have been returned to the desert floor

Author’s name: Annalea Beattie

Mission Status: Nominal

Sol Summary Activity: We began our day early and Clare, Mehnaz and I left for our EVA at 9.30 am, (#12 EVA to the left and right of the road at Marble Ritual and Pooh’s Corner). As part of Clare’s geoconservation research strategies (and because it is the right thing to do) we returned rock samples from the hab to three of their original sampling sites. This involved a bit of walking but we did complete this EVA well ahead of time. The weather was shaky, turned cold and looking like rain. As we had completed our goals we came home early to the smell of baking, to Dan’s tri-colour Mars Flag cake and Aditya’s French bread. Tough for us. In the afternoon the second EVA (# 13 EVA to the Dry Creek and Cowboy Corner) consisting of Rajvi, Aditya, Clare and Bharti, efficiently deposited all samples back to their original locations. After their return, our crew mostly wrote about their research. We are nearing the end of our time here and, as always when you are leaving somewhere great, there is work that you would still like to do. Late afternoon Daniel, Clare and I extracted DNA from a soil sample and some kind of uncertain yellow goo found in a puddle at Kissing Camel Ridge. As Crew Biologist, Daniel has freely shared his knowledge and his experience with all of us. He has included us in his work extracting DNA. Mehnaz and Rajvi completed their usual routine chores around the station and the Green Hab. Then Mehnaz interviewed both Rajvi and Aditya about their work. Report writing tonight, some cards maybe if we have time, pasta Napoli with fresh basil, Aditya’s French bread and a tomato and herb salad from the Green Hab for dinner.

Look Ahead Plan: We finish our simulation tomorrow at midday. In the morning we will complete all documentation, the End-of-Mission Report and then thoroughly clean the Science Dome. After lunch we will drive to the Valley of the Stars and to the box canyon at White Cliffs, to see at close range authigenic gypsum in the Sommerville.

Anomalies in work: None

Weather: 54F maximum today with blue skies then cloud cover then blue skies again, with wind gusts to 14 mph. There was a 1% chance of rain this afternoon, but this came to nothing. A mild dry day in the desert.

Crew Physical Status: Nominal

EVA: none

Reports to be filed: Sol Summary, Journalist Report, Green Hab Report, Operations Report, EVA report, EVA request.

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site